### v1

# aims-PAX: Parallel Active eXploration for the automated construction of Machine Learning Force Fields

Abstract
Recent advances in machine learning force fields (MLFF) have significantly extended the capabilities of atomistic simulations. This progress highlights the critical need for reliable reference datasets, accurate MLFFs, and, crucially, efficient active learning strategies to enable robust modeling of complex chemical and materials systems. Here, we introduce aims-PAX, an automated, multi-trajectory active learning framework that streamlines the development of MLFFs. Designed for both experts and newcomers, aims-PAX offers a modular, high-performance workflow that couples flexible sampling with scalable training across CPU and GPU architectures. Built on the widely adopted ab initio code FHI-aims, the framework seamlessly integrates with state-of-the-art ML models and supports pretraining using general-purpose (or "foundational") models for rapid deployment in diverse systems. We demonstrate the capabilities of aims-PAX on two challenging systems: a highly flexible peptide and bulk CsPbI3 perovskite. Across these cases, aims-PAX achieves a reduction of up to two orders of magnitude in the number of required reference calculations and enables over 20x speedup in AL cycle time through optimized resource utilization. This positions aims-PAX as a powerful and versatile platform for next-generation ML-driven atomistic simulations in both academic and industrial settings.
preprint: AIP/123-QED
I Introduction:
The success of Machine Learning Force Fields (MLFFs)Unke et al. (2021) has deeply transformed the field of molecular simulations. They are now the preferred method for simulating the dynamics of large systems, such as extensive perovskite unit cellsBaldwin et al. (2024) or solvated proteinsUnke et al. (2024), with quantum-chemical accuracy. While general-purpose (GP) (sometimes called “foundational”) modelsKabylda et al. (2025); Smith, Isayev, and Roitberg (2017); Kovács et al. (2025); Chen et al. (2019); Deng et al. (2023a); Choudhary and DeCost (2021); Merchant et al. (2023); Gasteiger, Becker, and Günnemann (2021) trained on large datasetsHoja et al. (2021); Eastman et al. (2023); Deng et al. (2023b); Barroso-Luque et al. (2024); Levine et al. (2025); Ganscha et al. (2025) are becoming more widespread, there remains a strong demand for high-quality data to fine-tune these models or to build new custom models for challenging applications.Novelli et al. (2024); Kaur et al. (2025); Radova et al. (2025)
The process of collecting representative high-quality datasets can be labor-intensive, requiring considerable manual effort and computational resources. To address these challenges, a common approach is to employ active learning (AL).Settles (2009); Ren et al. (2021) In AL, an uncertainty measure of a model’s prediction is used to select data points for labeling and inclusion in the training dataset. This approach enriches the training dataset with data points that represent a challenge for the current state of the model. In essence, the model autonomously determines which data to prioritize for training and which to disregard. Therefore, this procedure reduces human intervention and decreases the computational cost of model training by requiring only a small number of costly and slow reference ab initio calculations to reach an acceptable accuracy. In addition, it also improves the robustness of the MLFF by auto-detecting and correcting its possible failures during the training procedure.
AL has been successfully applied to a plethora of applications. For example, Young et al.Young et al. (2021) used active learning to iteratively improve a MLFF that was able to accurately simulate solvents and selected chemical reactions. In a study by Stark et al.Stark et al. (2023), an AL workflow leveraging clustering algorithms was used to model reactive hydrogen dynamics on copper surfaces. Furthermore, Mohanty et al.Mohanty et al. (2023) showed how AL was necessary to augment a dataset for efficiently training MLFFs for polymer dynamics and Kang et al.Kang et al. (2024a) highlighted how AL was crucial to model strongly anharmonic materials. Numerous other successful AL applications can be found in the literature. Zhang et al. (2019a); Smith et al. (2018); Zhang et al. (2023); Shambhawi, Csányi, and Lapkin (2021); Sivaraman et al. (2020); Kuryla et al. (2025); Erhard et al. (2024); Vandermause et al. (2024); Duschatko et al. (2024); Xie et al. (2023); Vandermause et al. (2022); Johansson et al. (2022); Xie et al. (2021); Matsumura et al. (2025); Gurlek et al. (2025) While AL is always beneficial in the data collection process, the automation degree of the procedure varies broadly. Often, AL is done manually or by users who develop tailored scripts for their specific problems. This situation results in the need for expert knowledge, such as selecting starting geometries, setting uncertainty thresholds, or deciding when to stop sampling. Additionally, employing collections of custom scripts instead of a defined workflow makes the process less accessible to new practitioners and less reproducible by other researchers. In recent years, the community has started addressing these challenges by offering various automated software solutions. For example, in the DFT codes such as the Vienna ab initio simulation package (VASP)Kresse and Hafner (1993); Kresse and Furthmüller (1996a, b), CASTEPClark et al. (2005); Stenczel et al. (2023) and the Amsterdam Modeling Suite (AMS)Rüger et al. (2025) different automated AL workflows are implemented. Next to AL methods directly integrated into quantum chemistry codes, there also exist separate software packages offering AL or automated simulation functionalities such as FLAREVandermause et al. (2020), CatFlowLiu et al. (2025), AlebrewZaverkin et al. (2024), PsiFlowVandenhaute et al. (2023), ALmoMDKang et al. (2024b), apaxSchäfer et al. or PALZhou et al. (2024). Although such tools have helped establish MLFFs and AL as a standard tool in molecular simulations, there is a potential for improvements that we address in this work, in particular with respect to the efficiency of configurational space exploration, hardware utilization and support for multisystem sampling.
We present aims-PAX, short for ab initio molecular simulation-Parallel Active eXploration, a flexible, fully automated, open-source software package for performing AL using a parallelized algorithm that enables efficient resource management. The current implementation is coupled to the FHI-aimsAbbott et al. (2025) program for DFT calculations and the MACEBatatia et al. (2022, 2025) architecture as an MLFF model. However, the algorithm itself is agnostic to both the MLFF architecture and the DFT code.
We showcase aims-PAX performance by applying it to a large, flexible peptide. In the case of this peptide, aims-PAX automatically produces accurate and stable MLFFs and provides a high-quality dataset with a minimal number of ab initio calculations. For the latter, we compare aims-PAX performance to a more traditional manual data collection strategy used in the TEA challenge.Poltavsky et al. (2025a, b) Fully automated, aims-PAX reduced the number of required DFT calculations by two orders of magnitude while achieving essentially the same results.
The structure of the article is the following. First, in the Methods section we provide a technical description of the aims-PAX workflow, as well as a thorough report of computational methods and their settings used. In the Results section, we report the application of aims-PAX to the peptide Ac-F-A5-K. In addition, the computational advantages of the parallelized AL algorithm are shown on the example of bulk CsPbI3 pervoskite. We conclude the article with a short summary and outlook for the future of aims-PAX.
II Methods
In this section, we describe the proposed aims-PAX AL workflow by first explaining how initial datasets and initial models can be generated efficiently through an integrated procedure that leverages GP-MLFFs. Afterwards, we explain the core of the proposed method, which consists of the parallelized AL algorithm. We describe the implementation details, multi-trajectory sampling, and the employed uncertainty measure, along with its adaptive threshold. We close the section with a comprehensive account of the computational methods used in the various studies.
II.1 Initial Dataset and Model Generation
A starting point for an AL procedure involves generating an initial ensemble of MLFFs, or a single MLFF, that simultaneously predicts the potential energy surface (PES) and associated uncertainties, capable of producing stable molecular dynamics within a limited region of the PES. We want to emphasize that this part of the workflow is not active in the sense that the model does not choose which points to include in the training. At this stage, the model is not yet sufficiently reliable to guide this selection process. Thus, data is generated using a sampling strategy, such as molecular dynamics (MD).
The described initial dataset generation (IDG) can be established using one of two approaches:
  1. 1.
    Short ab initio simulations can be run to generate molecular configurations along with their respective energies, forces etc.
  2. 2.
    A GP-MLFF can be used to produce physically plausible system geometries. These geometries are then recomputed using a reference ab initio method.
The second approach is generally preferable, as it helps decorrelate the geometries, making the IDG significantly more computationally efficient—by at least an order of magnitude. Also, we coupled the IDG with ParslBabuji et al. (2019), similar to what is done in PsiflowVandenhaute et al. (2023). This enables users to perform DFT calculations on sampled geometries across multiple nodes in parallel. Importantly, the GP model does not need to provide accurate energies or forces; it acts solely as a geometry generator in combination with MD simulations.
Both initialization strategies are implemented in aims-PAX, see Fig. 1b. Currently, the implementation includes the MACE-MP0Batatia et al. (2024) model, with additional models to be incorporated in the future, such as SO3LR Kabylda et al. (2025).
Once the dataset reaches a user-defined threshold in size, it is split into several equally sized subsets. These subsets are randomly selected from the full dataset, with one subset being assigned to each MLFF ensemble member. This ensures that each model is trained and validated on slightly different data. In addition to varying the training data, we introduce further diversity between ensemble members by using different random seeds for initializing model weights. Each MLFF is then trained on its assigned subset for a user-specified number of epochs.
Optionally, this procedure of generating data and training can be repeated multiple times. More precisely, after training, new structures are sampled and subsequently labeled, which is followed by more training steps. More details on this approach can be found in Section A.3.1.
During these cycles, the MLFFs are trained without reinitializing their weights. Instead, the existing weights are reused at every training step, and models are trained on their entire datasets to prevent catastrophic forgettingFrench (1999); McCloskey and Cohen (1989); McClelland, McNaughton, and O’Reilly (1995); Ratcliff (1990); Kumaran, Hassabis, and McClelland (2016). This continual learning (CL) approach Kirkpatrick et al. (2017); Aljundi (2019); Parisi et al. (2019) enables models to improve iteratively without retraining from scratch each time, reducing the number of required training steps. Crucially, at this stage, models are not trained to full convergence; instead, training is deliberately limited to a small number of epochs (typically five or fewer). The described early-stopping strategy avoids overfitting the models on the initial datasets and hinders them from getting stuck in local minima. This would make updating the models with new data during the AL significantly more difficult without reinitializing their weights.
The IDG is repeated until a user-defined stopping criterion is met. Possible stopping criteria include a maximum number of training epochs, a predefined training set size, or a target performance (e.g., force mean absolute error, MAE) on the validation set. The latter can be aligned with the overall AL workflow termination condition. For instance, the user may specify a target force MAE that should be achieved on the validation dataset by the end of the AL process. A scaling factor can be applied to this target MAE to define the stopping criterion for MLFF ensemble pretraining. At this stage, the goal is not to develop highly accurate MLFFs or exhaustive datasets but to obtain a robust MLFF ensemble capable of generating stable dynamics within an initial region of the PES, from which the main AL workflow can begin sampling the broader PES landscape.
II.2 Parallelized Active Learning
The AL phase involves sampling the configurational space of the target system using a pre-trained ensemble of MLFFs, which are employed for both sampling and uncertainty quantification. The latter is used together with a threshold that determines when a sampled structure is supposed to be labeled via a DFT calculation.
In the case of aims-PAX, each time the threshold is crossed and the DFT calculation has been performed, the new data is added to the training (or validation) set of all MLFFs. These are then updated in a CL scheme using a user-specified, ideally low, number of epochs similarly to the one employed in the IDG. For more details on the exact training strategy we refer to Section A.3.1.
While the algorithm proposed herein is, in principle, agnostic to the choice of uncertainty quantification method, we employ the query by committee (QBC)Seung, Opper, and Sompolinsky (1992); Schran, Brezina, and Marsalek (2020); Artrith and Behler (2012) approach due to its conceptual simplicity and widespread adoption. The integration of alternative uncertainty estimation techniques into our framework is straightforward and will be explored in future work.
In the QBC approach, an ensemble of independently trained ML models is used to produce a distribution of predicted outputs during inference. As described previously, diversity among ensemble members arises from differences in initial weight initialization seeds and distinct initial training datasets. The variance within the ensemble predictions serves as a way to quantify a model’s uncertainty. Specifically, we quantify uncertainty based on the variance of atomic force predictions, using the maximum per-atom force variance across the system, as defined in Eq. 1Zhang et al. (2019a),
$\delta_{n}=\underset{i}{\max}\sqrt{\frac{1}{3M}\sum^{M}_{j=1}\sum_{k\in{x,y,z}}\left(F_{nijk}-\bar{F}_{nik}\right)^{2}},$ (1)
where $\delta_{n}$ denotes the uncertainty associated with geometry $n$. The maximum is computed over all atoms $i$ in the system. The ensemble consists of $M$ models indexed by $j$, and the summation over $k$ spans the three spatial components $x$, $y$, and $z$. The term $F_{nijk}$ represents the $k$-th Cartesian component of the force on atom $i$ in system $n$ predicted by model $j$, while $\bar{F}_{nik}$ denotes the ensemble-averaged force component on atom $i$ in direction $k$.
For setting the uncertainty threshold, we adopt an approach analogous to the one implemented in VASPKresse and Hafner (1993); Kresse and Furthmüller (1996a, b), where a scaled moving average of the uncertainties is used in place of a fixed threshold. Specifically, the threshold at iteration $t$, denoted by $\delta_{t}$, is computed using Eq. 2,
$\delta_{t}=\frac{1+c_{x}}{N}\sum_{n=1}^{N}\delta_{n}.$ (2)
Here, $N$ represents the number of past uncertainty values included in the moving average, for which we follow definitions introduced in the VASP code and use a default window size of 400. The scaling factor $c_{x}$ allows the threshold to be adjusted: values $c_{x}<0$ tighten the threshold, while $c_{x}>0$ relax it. In our implementation, the default value is $c_{x}=0$. We also include the option to freeze the threshold after a user-specified training set size.
The primary advantage of this adaptive thresholding approach is that it eliminates the need for a fixed, user-defined uncertainty cutoff, which can vary between systems.Kulichenko et al. (2023) Since the moving average naturally decreases over time, some configurations will always exceed the threshold. As a result, the sampling frequency depends on the value of $c_{x}$: if set too high, very few points may be sampled; if too low, the method may oversample. Based on our experience and also reported for the MLFF training in the VASP code, values of $c_{x}\in\left[-0.1,\,0.1\right]$ serve as practical starting points.
To improve the efficiency and robustness of the active sampling, we adopt a multi-trajectory approach that has also been successfully applied in similar frameworks.Liu et al. (2025); Zhou et al. (2024); Matsumura et al. (2025) Herein multiple ML-driven simulations are executed in parallel, see Fig. 1c. These trajectories may differ in their sampling strategies, utilizing various thermostats, barostats, external conditions, or simulation schemes ranging from classical MD to enhanced sampling techniques such as metadynamics. Although only a subset of these options is currently implemented, additional methods are under development and will be integrated in future updates. We point out that the uncertainty threshold as defined in Eq. 2 is shared across all of these trajectories.
A key advantage of multi-trajectory sampling is its ability to decouple the generation of new configurations from the evaluation of high-uncertainty states. While MLFFs generate new candidate geometries, DFT calculations are performed in parallel on selected high-uncertainty configurations to enrich the reference dataset. These calculations are done using FHI-aimsAbbott et al. (2025) compiled as a library and interfaced through the Atomic Simulation InterfaceStishenko et al. (2023). The latter allows to run an instance of FHI-aims continuously, meaning that the DFT code does not have to be reinitialized before every calculation. This can remove significant overhead, which is especially valuable when handling smaller systems.
As mentioned earlier the training is done using a CL scheme, similar to the one used during pre-training, which allows MLFFs to be incrementally updated during sampling. Together with the parallel DFT calculations, this strategy also optimizes utilization of available computational resources (CPUs and GPUs), thereby enhancing the overall efficiency and throughput of the AL workflow.
Similarly to the IDG, ParslBabuji et al. (2019) can also be used here, allowing to distribute DFT calculations across multiple nodes in parallel. Additionally, the number of DFT workers can be adapted dynamically up to a user-defined maximum. Through the flexible allocation of resources, we ensure that no worker is idle or overloaded. This is particularly useful in AL as the demand for new data can change during the procedure. For example, there can be long sequences where an MLFF trajectory is certain about all encountered geometries or intervals where all trajectories lead to high uncertainty regions.
As with the IDG, the AL workflow proceeds until a user-defined stopping criterion is met. This may be based on the total training set size, performance on the validation set (e.g., force MAE), number of training epochs, or total number of MD steps. Once the stopping criterion is met, either the entire ensemble or only the best-performing ML model, selected based on validation error, is further trained to converge on the whole training set.
Figure 1: Overview of the aims-PAX workflow. (a) Required input files: The first file (control.in) follows FHI-aims conventionsBlum et al. (2009) and contain the DFT settings and the system’s geometry, initial geometries can either be inside a folder (geometry/) or, in case of a single geometry, in a file (geometry.in), the third file (mace.yaml) contains MACEBatatia et al. (2022, 2025) model parameters and the fourth (aims_PAX.yaml) is an aims-PAX-specific file containing the AL settings. For the AL workflow, folders containing the initial datasets (data/) and models (models/) are required. (b) Initial dataset generation (IDG): Geometries are sampled using either DFT or a GP model, with DFT providing labels in both cases. Sampling continues until a (* user specified) criterion is met. (c) Parallelized active learning: The AL workflow requires input files, existing data, and models, which can be provided by the IDG procedure. Sampling occurs over multiple trajectories, triggering DFT calculations when an uncertainty threshold is exceeded. GPU-based ML tasks (orange) and CPU-based DFT tasks (blue) can run in parallel. AL is continued until a (* user specified) stopping condition is met. (d) Output: Models and collected data produced during AL (and IDG).
II.3 Computational Details
All DFT calculations were performed using FHI aimsBlum et al. (2009) version 241114 compiled as a library and called through the python ASI package asi4pyStishenko et al. (2023) version 1.3.18 connected with ASEHjorth Larsen et al. (2017) version 3.23.0. For MACEBatatia et al. (2022, 2025) we used mace-torch version 0.3.9. with pytorchPaszke et al. (2019) version 2.3.1. For aims-PAX with PARSL, we used PARSL version 2024.12.16 and FHI aimsBlum et al. (2009) compiled as an executable. In this implementation, the ASEHjorth Larsen et al. (2017) calculator is used to perform DFT calculations.
N-acetylphenylalanyl-pentaalanyl-lysine (Ac-F-A5-K)
During AL, we employed the Perdew-Burke-Ernzerhof (PBE) functionalPerdew, Burke, and Ernzerhof (1996) with non-local many-body dispersion (nl-MBD)Hermann and Tkatchenko (2020) using the LIGHT species defaults for numerical settings and basis sets. Relativistic corrections were applied using the atomic ZORA approximation.van Lenthe, Snijders, and Baerends (1996) The total energy, eigenvalue, density, and force convergence criteria were set to $10^{-6}$ eV, $10^{-4}$ eV, $10^{-5}$ e/Å3, and $10^{-4}$ eV/Å, respectively. For recomputing the dataset to a higher level of theory, the PBE0Adamo, Cossi, and Barone (1999) functional with the nl-MBD dispersion correction and the INTERMEDIATE species defaults for basis sets and numerical settings was used, keeping the other settings fixed.
The serial version of aims-PAX was used for both IDG and AL. The former was performed by sampling 8 points for each member of an ensemble of 4 models with a stopping criterion of a maximum of 50 epochs. The structures were sampled using the small MACE-MP0 GP modelBatatia et al. (2024) by running MD in the NVT ensemble with the Langevin thermostatBerendsen et al. (1984) at 500 K with a timestep of 1 fs and a friction coefficient of $0.001\leavevmode\nobreak\ \mathrm{fs^{-1}}$. In order to decorrelate the data points, structures were picked every 20th MD step. Their energies and forces where then computed using DFT.
The AL workflow was run until a training set size of 500 structures was reached with 1:1 ratio for the validation set. During the AL, when new data was added to the training set, the models were trained for a total of 10 epochs on the updated dataset. The training was split into two steps, each involving 5 epochs. More precisely, this means that after 5 epochs are trained, the other running trajectories are propagated first before finishing with 5 more epochs of training. For more details on this see Section A.3.1. During AL, the structures were sampled using the same MD settings as in the IDG. The uncertainty threshold parameter $c_{x}$ (see Eq. 2) was set to the default value of 0.
The test sets for Ac-F-A5-K were created by running MD with the small MACE-OFFKovács et al. (2025) potential in the NVT ensemble at 300, 500, and 700 K using the LangevinBerendsen et al. (1984) thermostat with a friction coefficient of $0.001\leavevmode\nobreak\ \mathrm{fs^{-1}}$ and a time step of 1 fs for 1 ns. Every 100th geometry was selected and from the remaining points 1000 were selected by farthest point samplingEldar et al. (1997) using the mean, invariant atomic MACE-OFF descriptors. The chosen geometries were then recomputed using FHI aims with the same functional and settings used in the AL.
The MD simulations for assessing the stability of models were performed in the NVT ensemble at 300, 500, and 700 K using the Langevin thermostat with a friction coefficient of $0.001\leavevmode\nobreak\ \mathrm{fs^{-1}}$ and a time step of 1 fs for 1 ns. Throughout the simulation, the bond lengths were monitored, and if any of them exceeded 2 Å, the simulation was stopped.
Bulk Perovskite (CsPbI3 2x2x2)
During the AL, the PBE functional with the pairwise Tkatchenko-Scheffler dispersion correctionTkatchenko and Scheffler (2009) was used, employing the INTERMEDIATE species defaults for numerical settings and basis sets. Relativistic corrections were applied using the atomic ZORA approximation.van Lenthe, Snijders, and Baerends (1996) The total energy, eigenvalue, density, and force convergence criteria were set to $10^{-6}$ eV, $10^{-5}$ eV, $10^{-5}$ e/Å3, and $10^{-4}$ eV/Å, respectively. A Gaussian smearing of $0.05$ eV was applied to the orbital occupations. The calculations employed a parallel KS method with load balancing and local indexing enabled. A maximum of $300$ self-consistency iterations was allowed. The charge mixing parameter was set to $0.02$. The k grid was set to $1\times 1\times 1$. The lattice vectors were [17.23958, 0, 0], [0, 17.23958, 0], and [0, 0, 25.00256], all in Ångstrom.
The parallel version of aims-PAX employing Parsl was used for both IDG and AL. The former was performed by sampling 10 points for each member of an ensemble of 4 models with a stopping criterion of 50 epochs for the initial training. The structures were sampled using the small MACE-MP0 GP modelBatatia et al. (2024) by running NPT MD with the Nosé-Hoover thermostatEvans and Holian (1985) at 300 K and the Parinello-Rahman barostatParrinello and Rahman (1981) at 1 bar. The timestep was set to 1 fs and otherwise default ASE parameters were used. In order to decorrelate the sampled data points, structures were picked every 20th MD step. Their energies, forces and stress where then computed using DFT. The models were converged on the initial dataset before continuing with AL by training them until no improvements w.r.t. the validation set was achieved for 50 epochs
The AL workflow was run until a training set size of 100 structures was reached with 7:3 ratio for the validation set. The maximum epochs per trajectory are 10, and the intermediate epochs are 10. The structures were sampled using the same MD settings as described in the IDG above. The uncertainty threshold parameter $c_{x}$ (see Eq. 2) was set to 0.2 and the uncertainty itself was checked every 10th MD step.
Settings for MACE
The MACE architectures used during the aims-PAX runs are summarized in Table 1.
System
Parameter Ac-F-A5-K (small) Ac-F-A5-K (large) CsPbI3
Channels 32 256 64
Max degree $L_{\text{max}}$ 1 2 1
Cutoff [Å] 6 6 6
Radial Bessel functions 8 8 10
Message-passing layers 2 2 2
Correlation order 3 3 3
Radial MLP layers 3 3 3
Neurons per MLP layer 32 64 16
Activation function SiLU SiLU SiLU
Output Layer Irreps "128x0e" "16x0e" "16x0e"
Table 1: Architectural parameters of the MACE models used for the systems studied in this work.
For the training during AL with aims-PAX the follwing settings were used. The AMSGrad optimizer Reddi, Kale, and Kumar (2018) and a learning rate of 0.01 were utilized throughout. For the the IDG, the learning rate was decreased by 0.8 using the Reduce On Plateau scheduler with a patience of 5 and $\gamma=0.9993$. No learning rate scheduler was used during the AL. An exponential moving average of 0.99 for the model parameters and a gradient clipping of 10 were used. For the loss function a weighted mean square loss of energies and forces with weights 1 and 1000, respectively, was utilized. A batch size of 5 was used throughout. After the AL runs themselves, the best performing models of the respective ensembles were trained on the final training set until there was no improvement w.r.t. the validation set for 50 epochs.
For training the large model for Ac-F-A5-K (third column in Table 1) from scratch on the recomputed dataset, the same settings as described above for aims-PAX were used except the following changes. The energy and force weights of the loss function were set to 44 and 1000, respectively. The model was trained for 1000 epochs and after the 750 epochs the energy and force weights were swapped and the learning rate set to 0.001. The learning rate was decreased by 0.8 using the Reduce On Plateau scheduler with a patience of 256 and $\gamma=0.9993$.
Hardware
The benchmarks, the active learning and training were performed using a NVIDIA Ampere 40 GB HBM GPU and AMD EPYC Rome 7452 CPU. Recomputing the data on a higher level of theory and the DFT calculations through Parsl were done using AMD EPYC Rome 7H12. The training of MACE models from scratch and MD runs for Ac-F-A5-K were performed using a NVIDIA A100 80GB GPU.
III Results
To demonstrate the performance of aims-PAX, we use it to generate reference datasets and train MLFFs for a challenging system: a N-acetylphenylalanyl-pentaalanyl-lysine (Ac-F-A5-K) peptide. This system was selected because of its complexity and relevance to typical MLFF applications in biochemistry. Our results demonstrate that the proposed AL framework reduces the number of required reference evaluations by up to three orders of magnitude and substantially minimizing the necessary human effort.
III.1 The phase-space of a peptide: Ac-F-A5-K
We begin by applying our workflow to generate a robust MLFF for the Ac-F-A5-K peptide, a system comprising 100 atoms. This peptide exhibits multiple local minima explored during MD simulations under ambient conditions, which typically require numerous costly reference calculations when using conventional, non-AL methods. To assess the reliability of uncertainty estimates within our AL workflow, we trained an ensemble of models for the Ac-F-A5-K peptide using the aims-PAX framework, based on three parallel MD trajectories, as detailed in Section II.3. During the AL process, we also perform DFT reference calculations every 50 MD steps independently from the uncertainty selection criterion. Additionally, the actual prediction error and model uncertainty were evaluated at these points. Using this data we analyze the behavior of the uncertainty measure throughout the AL procedure without a bias towards high uncertainty states.
Figure 2 (a) shows the evolution of prediction error, uncertainty threshold, and training set size over the course of an AL run for each trajectory.
All three trajectories display consistent behavior: model uncertainty, the uncertainty threshold, and prediction error all decrease systematically as AL progresses. Notably, the temporal profiles of uncertainty and error follow similar trends, indicating a positive correlation between these quantities. To quantify this observation, we plot the uncertainty against the maximum atomic force error in Figure 2 (b), along with a linear regression fit, and compute the Pearson correlation coefficient. Across all trajectories, we observe a clear positive correlation between uncertainty and error, with only a limited number of outliers. This positive correlation is crucial, confirming that ensemble uncertainty can serve as an effective proxy for prediction error. Consequently, the AL algorithm selectively targets challenging configurations for high-fidelity DFT calculations while avoiding redundant sampling of trivial structures. Importantly, a perfect agreement between uncertainty and error is not required for practical applications; some errors may be missed in the early stages but captured at later AL steps as more diverse geometries are encountered.
Despite the widespread use of the QBC strategy for MLFFs, concerns have been raised regarding its reliability.Lu et al. (2023) Also, we have chosen a relatively small ensemble size of 4 and it has been reported that small ensembles result in biased estimators of uncertainties and other properties Imbalzano et al. (2021). However, it is not unusual, to use only a few ensemble members for AL in MLFFs.Zhang et al. (2019a); Stark et al. (2023); Kuryla et al. (2025); Erhard et al. (2024); Kang et al. (2024a) This is often done to reduce the computational expense of an AL procedure as increasing the number of ensemble members means more ML models have to be trained and evaluated. We opted for a small ensemble because, as shown above, the uncertainty measure is already a good approximation for the real error with four members. Thus, there is no need to incur greater computational cost by using more MLFF models in the ensemble.
To further probe the reliability of our approach, we analyze the evolution of the Pearson correlation coefficient between uncertainty and error throughout the AL process, see Figure 2 (c). In the initial 3k MD steps, the correlation exceeds 0.5 for all three trajectories. However, the correlation declines from 3k to 9k MD steps, even turning negative for the third trajectory (green). This degradation in uncertainty quality may be attributed to increasing overlap among the training sets of individual ensemble members as the AL progresses. As the models are exposed to similar data, they tend to converge on the same underlying potential energy surface, thereby reducing ensemble diversity. Nonetheless, the use of multiple trajectories helps alleviate this issue. A significant correlation for even a single trajectory can drive effective data acquisition, ensuring the continued efficacy of the overall AL scheme.
Another important aspect of the proposed AL workflow is the influence of multiple concurrent trajectories on the sampling process. To evaluate how model accuracy depends on the number of parallel trajectories used during AL, we conducted multiple aims-PAX runs with varying numbers of concurrent MD simulations. We performed three independent AL runs with different random seeds per setup for statistical reliability. For subsequent tests, we selected the best-performing model (based on validation set accuracy) from each of these three independent runs, resulting in three models per setup. The test sets were generated by performing $1$ ns of NVT MD at 300, 500, and 700 K using the MACE-OFF (small) potentialKovács et al. (2025). Representative structures were selected from these trajectories using farthest point sampling (FPS) based on ML-derived descriptors. Reference energies and forces were then computed at the chosen level of theory. Additional computational details are provided in Section II.3.
[…]
To investigate whether the use of CL during AL influences the performance of the resulting MLFFs, we retrained new models from scratch using the datasets acquired throughout the AL process. These models were evaluated using the same protocol described above for multiple trajectories. No deterioration in performance was observed for the models trained with CL compared to those trained from scratch. Detailed results are presented in the Appendix (Section A.1). These results confirm that the continuous learning paradigm offers a more computationally efficient alternative to repeated retraining without compromising the accuracy or robustness of the final MLFF models.
An essential requirement for MLFFs is the stability of the resulting MD simulations. Fu et al. (2023a). We performed four 1 ns-long NVT MD runs with each of the three models at 300, 500, and 700 K. This resulted in a total of 12 MD runs per number of trajectories used in the AL procedure and temperature. We define a simulation as stable if no covalent bond in the system exceeds 2 Å, a condition that is not expected to be violated at the temperatures considered. For more details, see Section II.3.
The results of all MD stability tests are summarized in Table 3. At 300 K, all MD trajectories remained stable, regardless of the number of trajectories used during the AL process. At 500 K, the temperature used during AL, nearly all MD runs were also stable, with only one exception: a single unstable simulation was observed for the model trained using eight parallel AL trajectories. In contrast, at the elevated temperature of 700 K, instability was observed in at least one MD run for every MLFF model tested. Specifically, for models trained with 8 and 32 trajectories, 7 out of 12 MD simulations were unstable. For the model obtained from a single-trajectory AL run, 3 simulations were unstable. Finally, for models trained with 4 and 16 trajectories, 1 and 2 simulations were unstable, respectively. These results align with the sampling strategy used in AL: since the training data were collected at 500 K, it is expected that MD simulations at or below this temperature (e.g., 300 K and 500 K) remain stable, as the MLFF is unlikely to encounter configurations outside its training domain. At 700 K, however, the MD trajectories explore more diverse and potentially unseen regions of configuration space, which can lead to instability due to extrapolation beyond the model’s training domain.
Overall, no clear trend emerges linking the stability of the MLFFs to the number of trajectories used during AL. This suggests that, for the current AL setup, model robustness in MD simulations is not significantly affected by the number of concurrent sampling trajectories. This behavior may be attributed to all trajectories using the same sampling protocol, potentially limiting exploration diversity. Future work will explore diverse sampling strategies across trajectories during AL to improve coverage of the potential energy surface and enhance model robustness under elevated temperatures and extreme simulation conditions.
Finally, the most reliable validation of an MLFF model lies in evaluating its performance in realistic application scenarios. Here, we assess the model’s ability to reproduce the Ramachandran plots from molecular dynamics simulations conducted under ambient conditions. The procedure follows that of the TEA 2023 Challenge benchmark Poltavsky et al. (2025b). We take the Ramachandran plots produced by the MACE model trained on the complete dataset in Ref. Poltavsky et al., 2025b as a reference. We recomputed the structures sampled by the aims-PAX workflow, using three parallel AL trajectories, at the same level of theory employed in the TEA 2023 Challenge (PBE0+MBDNL/intermediate). A new MACE model was then trained using the same architecture and hyperparameters as the reference study. For further details, see Section II.3. This recomputation was necessary because, during the AL phase, we employed a smaller MACE model trained on PBE+MBDNL/light to reduce computational costs. Such sampling-by-proxy strategies are commonly used in MLFF development Unke et al. (2021), and we demonstrate here how aims-PAX can efficiently generate high-quality, diverse datasets with minimal DFT overhead.
The retrained model was used to perform 12 independent $1$ ns NVT MD simulations at 300 K, each initialized from a different starting geometry, following the protocol of the TEA 2023 Challenge. The resulting trajectories were analyzed to extract dihedral angle distributions, which were then clustered following the methodology from Ref. Poltavsky et al., 2025b. The Ramachandran plots obtained from our model and the TEA reference model are shown in Fig. 2 (d). Both this work’s and the reference MACE models yield nearly identical cluster structures and populations for dihedral angles A and B, which correspond to the peptide backbone. Specifically, for angle A, a single dominant cluster is located at approximately ($\pi/4$, $\pi/2$) (blue), with a relative population of 100%. For angle B, two clusters appear in both models: one at ($\pi/4$, $\pi/2$) (blue) and another at ($\pi/2$, $\pi$) (green), with relative populations of 84% and 16%, respectively.
Minor differences are observed only in the dihedral angle C, which pertains to the peptide tail. Both models identify three clusters at similar angular positions, but relative populations differ slightly. For the blue cluster at ($\pi/2$, $\pi/2$), our model predicts a population of 64%, compared to 60% in the reference model. The orange cluster at ($\pm\pi$, 0) appears with a population of 7% in our model and 3% in the reference. The green cluster at ($\pm\pi$, $\pi/2$) is equally represented in both cases, with a population of 33%. The observed differences between MD results can likely be attributed to limited sampling statistics, as capturing slow conformational changes at the peptide tails may require simulations significantly longer than 12 ns.
Figure 2: Results from applying aims-PAX to the peptide Ac-F-A5-K: (a) Model uncertainty, actual maximum force error, uncertainty threshold and training set size as a function of MD steps throughout the AL procedure. (b) Actual maximum force error vs. model uncertainty with Pearson correlation coefficient over the whole AL workflow. A linear fit is shown as a guide to the eye. (c) Pearson correlation coefficient and training set size over multiple segments of the AL workflow for $n=3$ trajectories that were used for sampling. (d) Strucutre of Ac-F-A5-K including highlighting of relevant dihedral angles A,B and C. Ramachandran plot for said angles acquired with a model used in the TEA challengePoltavsky et al. (2025a, b) (left) and ours, acquired using aims-PAX (right). Relative populations of highlighted clustered are given in bold font (black) and the blue number in the bottom left corner of each plot indicates the percentage of configurations from the MD trajectories assigned to a cluster.Poltavsky et al. (2025a, b)
A crucial advantage of the proposed AL workflow is that our model was trained on only 500 reference structures, requiring a total of just 1,000 DFT calculations—including those performed during the AL process and the subsequent recomputation at a higher level of theory (excluding validation structures). In comparison, the reference model in Ref. Poltavsky et al., 2025b was trained on 4,000 structures generated from 100,000 DFT calculations, a process that also involved several months of manual effort. These results highlight the efficiency and scalability of the aims-PAX framework for the automated generation of high-quality training datasets. In particular, we demonstrate a reduced number of DFT evaluations by up to three orders of magnitude, achieved with minimal human intervention, while obtaining a final MACE model that delivers comparable predictive performance. Future developments, including the implementation of more reliable uncertainty quantification methods and diverse sampling techniques, are expected to strengthen further the advantages of the proposed automated AL workflow over traditional dataset generation and MLFF training approaches.
MAE Forces (meV/Å)
# Traj. 300 K 500 K 700 K
1 17.93 $\pm$ 2.28 27.07 $\pm$ 0.25 37.73 $\pm$ 0.84
4 16.60 $\pm$ 0.44 28.50 $\pm$ 0.61 37.47 $\pm$ 1.11
8 16.43 $\pm$ 0.67 28.73 $\pm$ 0.99 37.93 $\pm$ 1.36
16 17.20 $\pm$ 0.82 29.67 $\pm$ 0.95 39.70 $\pm$ 1.77
32 16.83 $\pm$ 0.38 29.77 $\pm$ 1.30 38.73 $\pm$ 1.36
Table 2: Mean absolute test errors of models acquired using aims-PAX with various number of trajectories used for sampling. Average and standard deviation over the best models from three separate aims-PAX runs.
# Stable MD Runs
# Traj. 300 K 500 K 700 K
1 12 12 9
4 12 12 11
8 12 11 7
16 12 12 10
32 12 12 7
Table 3: Number of stable MD simulations performed with models acquired using aims-PAX with various number of trajectories used for sampling at multiple temperatures. Three models were obtained from separate aims-PAX runs and four simulations were run for each. Stability was defined as no bonded atoms separated by more than 2 Å.
III.2 Computational Benefits of Parallelized Active Learning
In order to investigate the efficiency of the proposed parallel AL algorithm we choose to run aims-PAX in parallel and serial mode for the small peptide Ac-Ac-A3-NHMe (42 atoms) and the perovskite CsPbI3 (160 atoms in the unit cell). More detail regarding the exact settings for DFT and aims-PAX are described in Section II.3.
As the number of trajectories using in aims-PAX is integral part of the parallel algorithm, we ran the procedure with 4, 8, 16 and 32 trajectories for Ac-A3-NHMe. We used the completely serial version (the MLFF waits for DFT calculations which themselves are processed serially) and the CPU/GPU parallel version (the MLFF does not wait for DFT calculations but the latter wait for each other). The latter works through an MPI-based implementation, using ASIStishenko et al. (2023). The whole workflow has 1 GPU card as well as 1 CPU node with 128 cores available irrespective of how many trajectories are being used. The results of this study are shown in the appendix in Section A.2.
Through the implementation using Parsl we can easily distribute DFT calculations across multiple nodes dynamically. Therefore, the AL becomes CPU/GPU parallel, i.e., DFT calculations can run while the MLFF is being used, and CPU/CPU parallel, i.e., multiple DFT calculations can run in parallel. To investigate the advantage of this approach, we fix the number of trajectories to 32 and consider 1, 2, 4, 8, 16 and 32 CPU available nodes. The test is performed for the perovskite CsPbI3. We emphasize the number of available nodes here, as aims-PAX automatically scales up or down the number of workers (with one node per worker in our case) up to a user-defined maximum.
One DFT calculation, using the hardware and settings described in Section II for this perovskite takes around 20 user-minutes. Therefore, contrary to smaller systems where the bottleneck of the AL run can be the MLFF computation time, in the AL procedure of the perovskite the bottleneck will always be due to the DFT calculations. This is why we chose this system to investigate the scaling of aims-PAX to more workers. The results of the scaling test are shown in Fig. 3. Running aims-PAX with only 1 available node takes about 57 hrs of wall-clock runtime. By going to two nodes, the time spent is reduced to roughly 29 h, i.e. by a factor of nearly 2. Doubling the number of nodes to 4 and then 8, halves the run time to 14 h and 7 h, respectively.
The deviation from ideal scaling observed when using 16 and 32 nodes, as indicated by the decreasing slope in Fig. 3, stem from the fixed number of sampling trajectories, which is 32. The likelihood that all 32 (or even 16) trajectories require halting simultaneously, and thus trigger concurrent DFT evaluations, is low. As a result, the computational resources on all nodes are not fully utilized at all times. Overall, the use of the parallel aims-PAX implementation can reduce MLFF creation time by a large factor for systems requiring computationally demanding reference labeling, while efficiently utilizing both GPU and CPU resources.
Figure 3: Wall-clock runtime in hours as a function of the number of available CPU nodes for aims-PAX using Parsl applied to the pervoskite CsPbI3.
IV Conclusion
In this paper, we introduced aims-PAX, a flexible, fully automated, open-source software package for performing AL using a parallelized algorithm that ensures efficient resource management. The key advantages of the proposed workflow include: minimal human intervention; the use of state-of-the-art GP-MLFF models for initial dataset generation and pretraining; and a parallel workload manager that effectively utilizes all available computational resources. The AL process is distributed across multiple sampling runs. Reference DFT calculations can proceed independently whenever a member of the MLFF ensemble exhibits uncertainty regarding a newly encountered geometry.
The performance of aims-PAX is demonstrated on two diverse and challenging systems: Ac-F-A5-K, a highly flexible peptide; and the bulk perovskite CsPbI3. For the Ac-F-A5-K aims-PAX reduces the number of required DFT calculations by two orders of magnitude compared to traditional sampling approaches, providing robust and accurate MLFF suitable for running long MD simulations. Finally, using the example of bulk CsPbI3 perovskite, we demonstrate the advantage of parallel multi-trajectory sampling, reducing the AL time by an order of magnitude for systems requiring demanding DFT calculations. The presented aims-PAX software package can accomplish all these out of the box with minimal human effort.
We are convinced that the model- and ab initio method-agnostic nature of aims-PAX will be the basis of fruitful open-source collaboration. We envision multiple technical and practical avenues to explore. Optimizing the algorithm, e.g. by parallelizing across multiple GPU devices, will improve the efficiency and scalability of aims-PAX even further. Also, we plan to add support for automated fine-tuning of GP models as well as sampling across multiple chemical species. The latter will enable the automated creation of diverse, information-rich datasets and, together with the fine-tuning capabilities, pave the way to expand the chemical space of pre-trained MLFFs. Also, while we showcased aims-PAX with selected examples in this study, a central focus of current work is its application to complex and extended systems.
Thus, aims-PAX stands as a powerful, adaptable framework, enabling new frontiers in MLFF development and their ability to address complex chemical challenges.
Acknowledgements.
T.H. thanks Sander Vandenhaute for his time and valuable suggestions concerning the Psiflow and Parsl packages. He also thanks Iryna Knysh for her support in finding bugs inside the code.
We acknowledge financial support from the Luxembourg National Research (FNR) under the AFR project 17932705. The simulations were performed on the HPC facilities of the University of Luxembourg, (see hpc.uni.lu), the Luxembourg national supercomputer MeluXina, the computing resources at the Max Planck Institute for the Stucture and Dynamics of Matter in Hamburg and at the MPCDF. The authors gratefully acknowledge the LuxProvide teams for their expert support. S.S. acknowledges support from the UFAST International Max Planck Research School.
Data Availability Statement
The data and models used for the results in this study can be found on: 10.5281/zenodo.16893060
The code of aims PAX is available under the Github repository: github.com/tohenkes/aims-PAX.
Appendix A Appendixes
A.1 Training models from scratch
In order to assess the difference between continuously training the models during AL and training models from scratch afterwards, we trained 3 models with different seeds on the dataset of Ac-F-A5-K, acquired through the aims-PAX run with 4 trajectories as described in Section II.3. All settings for training, testing, MD simulations and the model architecture were kept the same. The results for the accuracy and stability are shown in Table 4. Comparing with results in Tables 2 and 3, there is no meaningful difference between training from scratch and using continual learning (CL) observable for accuracy and stability. Given that CL is computational more efficient, it is the mode of action used in aims-PAX.
Temperature MAE Forces (meV/Å) # Stable MD Runs
300 K 17.53 $\pm$ 0.90 12
500 K 28.03 $\pm$ 1.03 12
700 K 37.77 $\pm$ 1.25 10
Table 4: Mean absolute test errors (mean $\pm$ standard deviation) and number of stable MD simulations at various temperatures for models trained from scratch on data acquired via an aims-PAX run for Ac-F-A5-K. MAE values are averaged over three models with different seeds. Each model was used to generate 4 MD simulations, for a total of 12 per temperature. Stability was defined as no bonded atoms separated by more than 2 Å.
A.2 Speedup of CPU/GPU parallel aims-PAX version
Figure 4: Speedup of the parallel version of aims-PAX over the serial version applied to Ac-Ala3-NHMe running on a single CPU node with 128 cores and 1 GPU card.
The results regarding the speedup using the CPU/GPU parallel aims-PAX version are summarized in Fig. 4. We found that using 4 or more trajectories results in a speedup of 15 %. The largest speed-up can be observed with 8 trajectories with 18 %. For 16 and 32 the speedup is slightly lower than for 8 trajectories at 16 % for both runs. The speed-up observed when enabling more trajectories is caused by DFT calculations that run concurrently with the sampling of new points and training. Furthermore, a slight speed-up is observed between 4 and 8 trajectories, likely due to the reduced probability of no trajectory being propagated at a given time. Because trajectories are stopped when the uncertainty threshold is exceeded, it is more likely that all trajectories will stop if there are fewer of them, thus halting all AL progress.
Computational details
The same DFT settings used for Ac-F-A5-K were used for Ac-A3-NHMe during AL. The only exception is that no dispersion correction was applied. For generating the initial data set using aims-PAX the same settings as for Ac-F-A5-K were used.
For the AL workflow with aims-PAX, the same settings were used as for Ac-F-A5-K, except that the parallel aims-PAX version was used and the procedure stopped when the training set size of 200 was reached. Also, the same MACE architecture as described in Table 1 under Ac-F-A5-K (small) was used.
A.3 Technical Details of aims-PAX
Here we provide some technical details of the inner workings of aims-PAX for interested readers. We focus on the training during AL, explain how data as well as failed SCF convergence are handled and highlight slight differences between the order of operations in the serial and parallel AL algorithm.
A.3.1 Training during Initial Dataset Generation and Active Learning
During the initial dataset generation, the user decides how many points are selected for each ensemble member during each sampling step. For example, the user specifies that 5 structures per member are to be selected. aims-PAX then runs the sampling algorithm, and once 5 points were picked from the trajectory for each member (and labels are computed), the models are trained for a user specified number of epochs, namely intermediate_epochs. This process is repeated multiple times i.e. running sampling, picking points, labeling and training. The rationale behind this is, that if the user wants to have a specific accuracy of the models before running AL, aims-PAX makes sure that not too many structures are sampled and DFT calculations are performed.
During the active learning procedure each trajectory is associated with a state. Technically speaking, a loop is performed over all trajectories and depending on their state, different actions are performed. At the beginning of AL, all trajectories have the state running, which means the sampling algorithm is performed. Once a point is picked for labeling, the state of this specific trajectory is set to waiting until the DFT calculation is done and the results were received. This then changes the state to training and the user specified number of training epochs are performed, these are called intermediate_epochs_al. Afterwards, aims-PAX continues the loop over the trajectories. Only once a maximum, user-specified number of epochs, max_epochs_worker, is reached, the trajectory’s state swtiches back to running. This is done to enable other trajectories to continue sampling, potentially triggering new DFT calculations, which can then run while the models are trained. In addition, this means that the trajectories are always propagated with continuously updated model parameters.
A.3.2 Resetting the Optimizer during Active Learning
While training during AL, the weights are not reset when a new point is added. We have seen that repeatedly using the updated weights can result in the model being stuck in a minimum. We found it advantageous in this case to reset the optimizer state if the model is not improved after max_epochs_worker epochs (see Section A.3.1). This deletes the history of the adaptive optimizer (e.g. Adam or AMSGrad), resulting in a larger learning rate which helps the model to leave the local minimum.
A.3.3 Handling of new data points during Active Learning
Once a new point is selected and labeled during AL, it is either added in the training or validation set. To which dataset the new point contributes depends on a user specified ratio, that is kept consistent, e.g. 0.5, which means points are added to both sets alternately. In contrast to the IDG, both datasets are shared across models (except for the initial starting points that are present before the AL).
A.3.4 Handling of Failed SCF Convergence
During the IDG, if a DFT computation does not converge the geometry is discarded from the dataset. The procedure then just continues until any stopping criterion is met. This is especially relevant when using the GP model. However, we have not noticed any instances where SCF convergence could not be achieved for a geometry generated by the GP model for our systems.
In the case of AL, points where the SCF cycles do not converge are also discarded. On the trajectory where this is the case, a checkpoint geometry is loaded. This checkpoint is updated each time a selected structure is successfully labeled using DFT and the data is added to the training set. This ensures, that if the checkpoint is loaded, the MD continues from a geometry that is known to the MLFF.
A.3.5 Operational Differences: Serial vs. Parallel Version
While the overall AL workflow of aims-PAX is the same for its serial and parallel versions, there are slight differences that we want to point out. In the case of the serial procedure, the sampling and training of the ML models is halted if DFT calculations are performed. Afterwards, the model parameters are updated on the new data. Practically this results in all trajectories being propagated with the new information. For the parallel version, other trajectories can be propagated during the DFT calculation, meaning that sampling is done without the information of the current DFT calculation. While it can mean that potentially redundant points are sampled, the computational benefit, and thus possibility of scaling up the workflow, outweighs this inefficiency.